OpenStudy (anonymous):

Algebra 1- One Variable Equations Question: 3(x-1)+5x=6(x-2)-3 Okay, I know how to get through most of this equation, but I am finding it difficult to know what I should or shouldn't be subtracting/ adding to the equation to get the variable on one side. If any one could help, that would be greatly appreciated. I am not one of those students looking for just the answer, I know the answer from my text book. Sense this is a practice problem I would like to understand it fully. Thank you for your time, hope you can help! :) Thanks

OpenStudy (dtan5457):

What did you try? 3(x-1)+5x=6(x-2)-3 distribute both parenthesis first.

OpenStudy (anonymous):

I tried : 3x - 3 + 5x = 6x - 12 -3, I know you are suppose to get the variable to one side, but every time I try deducting or adding it doesn't come out to the answer of x = -6

OpenStudy (dtan5457):

3(x-1)+5x=6(x-2)-3 3x-3+5x=6x-12-3 combine like terms 8x-3=6x-15 subtract the 6x 2x-3=-15 add 3 2x=-12 x=-6

OpenStudy (anonymous):

Oh, wait now I understand you grouped the like terms: (3x + 5x) - 3 to get 8x - 3. Thank you; I was so focused on tying to decrease the amount of terms on one side, I totally forgot about needing to group them as well.

OpenStudy (dtan5457):

No problem. You seem eager to learn, I like that very much. Feel free to medal and fan, and ask me any more questions you may have :)

OpenStudy (anonymous):

Thanks. I fanned you, but I'm new to openstudy.com. So, please excuse my ignorance, how do you give someone a medal?

OpenStudy (dtan5457):

It's completely your choice when it comes to use, but just click "best response" to whatever answer helped you the most.

OpenStudy (triciaal):

@GhostOrchid welcome to openstudy!! Appreciate that you want to learn the material. All the best in your studies.

OpenStudy (anonymous):

Ah, okay. Well, hope you have a day/night!! As well, next time I will most defiantly be asking for you help if I get stuck.

OpenStudy (dtan5457):

Yes. The more eager you want to learn, the more eager people will want to help you. Personally, I refuse to help direct-answer seekers.

OpenStudy (anonymous):

dtan5457: No, I completely agree. If they don't want to help themselves, then why should you or any one else have to give up their own time to dedicate it to someone who doesn't care. But, honestly if my teacher wasn't off for the weekend I'd be calling lol. But, I don't think she appreciates my horrible math skills at times, so i try to figure it out on my own most of the time. But sadly, Math is fun.com, and khan academy. Only helped so far. :/ That's why I consider trying openstudy.com.

OpenStudy (dtan5457):

That's awesome, keep using this site with that mindset I and guarantee you will be able to help others and gain knowledge of your own. It's a win-win situation. I'll see you around the site.